Sol Summary – Oct 31st


Summary Title: Halloween Heliograph
Author’s name: Guy Murphy

Mission Status: Active.

Sol Activity Summary: EVA 4 was undertaken this morning to the chosen sample collection site for the micrometeorite study. During the EVA a heliograph was experimentally used to signal the hab from the field. Baseline micro/macro nutrients have been inputted into the food stock database. We completed tying the remaining untied tunnel tarpaulin eye loops to all the tunnels. 3 cooked meals prepared, including a tuna risotto with pesto sauce for dinner with an apple crumble desert. General hab housework duties were undertaken.

Look Ahead Plan: The EVA program will continue for the remainder of the week subject to favourable weather conditions.

Anomalies in work: None.

Weather: After a very cold night, it was another clear, sunny day with no precipitation.

Crew Physical Status: Crew in good health.

EVA: Successfully undertaken, see EVA report.

Reports to be filed: Sol Summary, EVA Report, Journalists Report, Science Report, Operations Reports, Photos.

Support Requested: We are still awaiting the delivery of food supplies, the absence of which is beginning to constrain our diet. Several plastic funnels to assist with the reuse of cooking water and several tea towels are also requested.

Science Report – Oct 31st

Crew 214 Science Report 31 October 2019
SOL: 4
Crew 214 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler

Science Operations: The second of the scientific EVAs was scheduled for the morning work period. This involved driving to site #2 at Tank Wash and establishing a 100 square metre (100 1m x 1m plots) on a relatively undisturbed natural surface. During the 2 hour EVA, the outline of the grid was established and three east-west delineators positioned. The task is incomplete.

A secondary experiment involved testing the magnetism of material brought to the surface at an adjacent anthill. Though it was hypothesized that this material might contaminate the sample plots; there was zero evidence of magnetic material at the surface.

A tertiary experiment was conducted at the Cactus Road #1 site before the return to the hab. With line of sight communication possible, a heliograph was used to demonstrate an emergency communications method should voice contact fail. The demonstration was successful and video records collected.

Tabulation of the baseline food stock macro and micronutrients for the waste study was completed.

Closed Loop Food Waste Study: The day’s waste for Sol 4 is: 229 grams; consisting of 891calories, 15g fat, 9g saturated fat, 1533mg sodium, 126g carbohydrates, 8g fibre, 4g sugar, 42g protein. Waste was significantly larger due to bread production errors. Otherwise, less waste is generally produced at breakfast, with almost all occurring at lunch or dinner, depending on the type of food cooked. The majority is preparation waste with no spoilage waste to date and a small amount of plate waste.

Additional Activities: Reinforcing the coverings for the tunnels continued throughout the day as time became available.

EVA Report – Oct 31st

Crew 214 EVA Report 31-Oct-2019
EVA # 4
Author: Andrew Wheeler

Purpose of EVA: Geological Sampling

Start time: 1000
End time: 1200

Narrative: The second EVA for science saw us traversing to the Tank Wash location #2 at 518477E 4253767N where we commenced setting out the 100 square metre (100 1m x 1m plots) sampling grid.

This location appears to be a relatively undisturbed natural surface that hasn’t undergone recent erosion or weathering. It is sufficiently away from the hab that metallic residue won’t contaminate the surface and appears to be upwind of vehicular and pedestrian activity.

A secondary experiment was a test of material brought to the surface by ants at a nest adjacent to the sampling grid. It was hypothesized that the magnetic material of a terrestrial origin may be present and would contaminate the micrometeorite sample material. Using a small neodymium magnet, it was observed that there was zero magnetic material located on the anthill.

A tertiary experiment was the use of a heliograph as an emergency communications method. There is a line of sight from the Cactus Road #1 location at 519570E 4251974N to the hab. With radio communications established and the hab personnel alerted, the heliograph operated as expected and was observed and recorded at the hab.

Destination: Drove from the hab along Cow Dung Road to Tank Wash and then returned to Galileo Road and the junction it has with Cactus Road.

Coordinates (use UTM NAD27 CONUS): 518477E 4253767N

Participants: Andrew Wheeler, Sandy Dance

Road(s) and routes per MDRS Map: Cow Dung Road to Tank Wash

Journalist Report – Oct 31st

Crew 214 Journalist Report
Thursday 31st Oct Sol 4
by Sandy Dance

A bit of a focus shift this morning when we realised we were using more water than we should be. So we sat around the table discussing how to reduce our water consumption: washing hands with sanitizer, washing dishes less often, keeping the vegetable rehydration water for future uses, washing ourselves even less often (we’re all in this together, so hopefully won’t notice!).

Andrew and I went out on an EVA this morning to install the string grid for the micrometeorite experiment in the location selected yesterday. This was a lot of fun, sticking satay sticks in the soft ground at the 4 corners of the grid, running 4 pieces of 10 metre string around the square, then the sticks at each 1 metre step around it. Fun until one skein of string decided to ravel into a hopeless knot, which took ages to unravel, in gloves. A lesson for Mars, find a way of controlling balls of string.

Anyway, it was a great success, we completed a good percentage of the grid as we expected, and were able to move onto the next exercise: running a magnet over a local anthill to see if the local rocks were magnetic. They weren’t, which should help the integrity of the micrometeorite experiment. We then moved to another location which had a line of sight back to the Hab and successfully sent a heliograph signal: a mirror flash.

Handy if all other forms of comms fail, if a little impractical.

The afternoon was spent in kick-back mode after our exertions, albeit including a bout of tying down more eyelets in the tarpaulins covering the tunnels.

Today David returned to the trailer outside the hub (or blasted off in the returning shuttle); it is a shame to lose my fellow vegetarian and his benign presence.

Operations Report – Oct 31st

Crew 214 Operations Report 31-Oct-2019

SOL: 4

Name of person filing report: David Mateus

Non-nominal systems: Opportunity, suit 10, generator

Notes on non-nominal systems: Oppy brakes do not work, Suit 10 is not charging and the generator stopped working in the night

Generator: run, Stopped running in the night

Hours run: N/A

From what time last night: 17:00

To what time this morning: 7:30

List any additional daytime hours when the generator was run: none

Solar— SOC % (Before generator is run at night): 77%

Diesel Reading –72 %

Station Propane Reading – 85 %

Ethanol Free Gasoline – N/A

Water (loft tank): 30 gallons

Water Meter: 1454950 units

Water (static tank) – 301 gallons

Static to Loft Pump used – No

Water in GreenHab – 150 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: It is not in the hab

Hours: –

Beginning charge: –

Ending charge: –

Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: –

Beginning charge: –

Ending charge: –

Currently charging: –

Spirit rover used: No

Hours: (before EVA): 106.1 h

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging):100%

Currently charging: Yes

Opportunity rover used: No

Hours: 65.5 h

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 112.5

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging):52%

Currently charging: Yes

Notes on rovers: No

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) No

Reason for use: None

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: Assigned to the director

HabCar used and why, where? Used by assistant director to go to town

CrewCar used and why, where? No used

General notes and comments: None

Summary of internet: functional

Summary of suits and radios: used for EVA, Suit 10 is not charging and we need batteries for 6 radios

Summary of Hab operations: All systems are nominal

Summary of GreenHab operations: No used

Summary of ScienceDome operations: All systems are nominal

Summary of RAM operations: No used

Summary of any observatory issues: No used

Summary of health and safety issues: All the crew are healthy

Questions, concerns, and requests to Mission Support: We need new batteries for six radios.

Final Schedule for the 2019-2020 Field Season

This one can be released to the public:

Final Schedule for the 2019-2020  field season:

Oct 17-20, 2019—Mars Society Conference, Los Angeles

Oct 22-25— IKEA Work Party

Crew 214—Oct 26- Nov10— MSA ExBoomerang

Crew 215—Nov 9-24, 2019— MSA ExBoomerang

Crew 216—Nov 23—Dec 8, 2019— A.R.E.S. (Mars Society crew)

Crew 217—Dec 7-14— Spaceward Bound Utah (Utah NASA Space Consortium Grant)

Dec 14-22— Holiday Break

Crew 218—Dec 21, 2019- Jan 5, 2020— Purdue

Crew 219—Jan 4-19 — ARES  (Florida Tech)

Crew 220—Jan 18-Feb 2— Mars Academy

Crew 221—Feb 1-16— Earth 2 Mars

Crew 222—Feb 15-Mar 1 —Supaero I

Crew 223—Feb 29-Mar 15— Supaero II

Crew 224—Mar 14-21—Wilderness Medical Society class

Crew 225—Mar 22-29—University of Colorado 2 Mars class

Crew 226—Mar 28-Apr 12–  Colombia

Crew 227—Apr 11-26—UCL

Crew 228—Apr 25-May 10—Mars Society II

Crew 229—May 9-24—Work Party

May 28-30– University Rover Challenge

Crew 230– June 1-8, 2020— Medical Makers

Crew 231—June 8-15—Martian Biology II

June 22-28—Astronomy Team Summer Residency

Copyright © The Mars Society. All rights reserved. | Main Site